Solve the following system by substitution
4x-12y=5
-x+3y=-1

Answer by jessbear13(1)

-x = -1 - 3y
x = 1 + 3y
4*(1+3y) - 12y = 5
4 + 12y - 12y = 5
4 = 5 therefore no solution